OpenStudy (anonymous):

Complete the square: 3x^2+7x-12 becomes x^2+7/4x=4 T or F?

OpenStudy (mathstudent55):

\( 3x^2+7x-12 = 0 \) First you must divide both sides by 3. you need \(x^2\) by itself.

OpenStudy (anonymous):

first , coefficient of x^2 must be 1,so divide by 3

OpenStudy (mathstudent55):

\( \dfrac{3}{3}x^2+\dfrac{7}{3} x -\dfrac{12}{3} = 0 \)

OpenStudy (anonymous):

x^2+7/3x+4=0?

OpenStudy (anonymous):

so it;s fasle?

OpenStudy (mathstudent55):

Good. \( x^2+\dfrac{7}{3} x - 4 = 0 \) \( x^2+\dfrac{7}{3} x = 4 \)

OpenStudy (mathstudent55):

False. You are correct.

OpenStudy (anonymous):

Yay! Thank you!